# Disentangling Vortex-driven Dynamics in the Solar Atmosphere Using Information Theory

### v1

@@ 2.5 Plasma heating
To identify the key drivers of temperature evolution within the vortex, we computed the NMI between plasma temperature and a selection of physical variables linked to heating processes or thermodynamic coupling. Figure 8 presents the temporal and vertical variation of these dependencies within the vortex core. The initial enhanced coupling between , and temperature is most pronounced during the vortex development and in the lower atmosphere, suggesting that the initial plasma cooling is driven by the jet transporting the vortex upwards and the initial flux tube expansion. As expected, plasma temperature initially shows strong statistical coupling with both density and pressure, consistent with the ideal gas law under quasi-static conditions. However, during the vortex phase, the coupling with plasma density significantly weakens above 1 Mm, indicating that the temperature evolution is no longer primarily governed by adiabatic compression or expansion. Instead, this points to a breakdown of ideal thermodynamic behavior due to non-adiabatic processes such as current dissipation, wave heating, and magnetically induced flows. The reduced NMI between temperature and density/pressure reflects the growing influence of magnetic forces and energy injection mechanisms in the upper atmosphere.
@@
Although vortex-driven expansion initially leads to adiabatic cooling, this effect is short-lived. Within 10 time steps, the plasma temperature begins to rise, indicating that additional heating mechanisms rapidly become dominant. The sustained NMI between temperature and magnetic shear throughout the vortex lifetime suggests that current structures play a consistent role in modulating plasma heating. This coupling is particularly notable given that the magnetic shear–temperature association emerges primarily during the vortex phase, vanishing before and after. This timing supports the interpretation that currents, associated with twisted and evolving magnetic fields, contribute to localized heating within the vortex core. The temporal alignment between this temperature rise and elevated magnetic shear points toward current dissipation as a likely source of energy input. Although this dissipation is numerical in origin, because of the lack of explicit resistivity in the simulation, it still reflects the physical tendency for currents to form and dissipate in regions of strong shear. This supports the interpretation that vortex-driven magnetic structuring can channel energy into thermal heating, especially under the low plasma- conditions of the chromosphere and transition region.
3 Discussions and Conclusion
@@
We examine two large magnetic flux tubes: one that hosts a strong solar tornado with a coherent vortex flow and another characterized only by shear flows and transient swirls. The vortex flow first emerges in the upper photosphere and lower chromosphere. From there, it expands both upward and downward, forming a long-lived column of rotating plasma. It persists for approximately 5–10 minutes, with a horizontal radius of over 1 Mm in the upper atmosphere. Our analysis focuses on its evolution from the photosphere through the upper atmosphere, within the height range of 0-4 Mm. Throughout the full height of the vortex, the plasma- steadily decreases as the structure develops, reflecting a growing dominance of magnetic forces over gas pressure. Interestingly, the magnetic field intensifies at the vortex onset and remains strong even after the large-scale structure dissolves, pointing to lasting magnetic reconfiguration as reported by Díaz-Castillo, S. M. et al. (2024). Aside from a brief initial cooling, the vortex phase is marked by a persistent temperature increase, highlighting the efficiency of heating mechanisms operating throughout its lifetime.
@@
Using information-theoretic diagnostics, SE and NMI, we quantified how vortex evolution alters MHD interactions. SE captures changes in the spatial complexity of individual variables, while NMI highlights how their couplings strengthen or weaken over time and height. The magnetic field variables illustrate this complementarity. SE shows that the entropy of the vertical magnetic field rises during vortex onset in the upper atmosphere, indicating the development of stronger spatial gradients and irregular distributions of the vertical magnetic field. As the vortex matures, entropy decreases, reflecting the emergence of a more ordered, collimated structure. By contrast, magnetic shear entropy begins low, rises sharply through vortex development as currents intensify, and falls again in the decay phase, consistent with current dissipation. Cross-comparison with NMI maps reveals that when magnetic shear entropy peaks, temperature becomes strongly coupled to shear, indicating that thermal evolution is linked to the buildup and subsequent dissipation of current-related complexity.
@@
The thermodynamic variables also show clear stratification. In the lower atmosphere, SE indicates low entropy in density but high entropy in pressure, reflecting relatively uniform mass loading alongside strong compressive variability. NMI confirms that temperature remains persistently coupled to horizontal compression () throughout the vortex lifetime. At the same time, temperature also shows strong coupling with vorticity, which itself contributes to velocity gradient. This dual dependence suggests that both compressive processes and viscous dissipation contribute to heating in the lower layers. In the upper atmosphere, the entropy patterns reverse: density entropy rises (stronger gradients and irregular distributions) while pressure entropy decreases (more ordered stratification). The coupling of temperature and density diminish considerably in the upper part of the domain, indicating departing from the ideal gas law. NMI reveals that temperature coupling with vorticity persists here, implying a continuing role for viscous heating, but now temperature also couples strongly to magnetic shear. This indicates that, higher up, magnetic structuring and current dissipation become additional drivers of thermal evolution alongside viscous effects. This confirms that vortices enhance the formation of both velocity and magnetic field gradients, creating favorable conditions for energy dissipation, consistent with previous studies (Yadav et al., 2021; Silva et al., 2024a, b). Our analysis extends this picture by showing that plasma temperature is statistically coupled to these gradients, demonstrating that the dissipated energy is effectively converted into heating within the vortex. Although the simulation does not include explicit resistivity, the hyperdiffusivity terms act on sharp gradients in the magnetic and velocity fields. In practice, this mimics the effect of current and viscous dissipation, meaning that regions where the vortex generates strong shear or currents correspond to sites of effective energy release. Thus, the observed thermal signatures can be interpreted as an analogue of vortex-driven heating in the solar atmosphere.
@@
The coupling is not one-sided: plasma variables also influence vorticity. Elevated temperature can enhance pressure gradients, which in turn drive flows that sustain or modify vortical motions. In the lower atmosphere, where pressure entropy is high, pressure gradients are likely to reinforce rotational shear and help maintain vortex coherence. In the upper atmosphere, strong magnetic shear can exert Lorentz forces that realign flows, thereby shaping or even destabilizing the vortex. The observed NMI links therefore suggest a bidirectional relationship: vorticity drives gradients that heat the plasma, while thermal and magnetic structuring provide feedback that regulates the strength, persistence, and eventual decay of the vortex as reported by (Shelyag_2012; Silva et al., 2024a)
The joint use of NMI and SE demonstrates that information theory provides a robust, flexible framework for probing nonlinear plasma interactions, even in data-limited settings. In particular, IT offers a pathway to assess causal dynamics without requiring full knowledge of the underlying governing equations. One of the major advantages of MI in plasma physics is that it bypasses some of the practical difficulties associated with working directly with the MHD equations. While the governing equations are complete in principle, in practice, they are highly nonlinear, coupled, and often intractable when one seeks to isolate the contribution of a specific process. Deriving the precise contribution of dynamic processes to plasma heating directly from the MHD equations is highly nontrivial. To even express temperature evolution analytically, one must assume an ideal gas law and a single-fluid approximation. Such assumptions break down in partially ionized, radiatively active, and magnetically structured regions like the solar chromosphere and in the vortex itself, as showed by our results. The energy equation includes complex terms such as anisotropic thermal conduction, optically thick radiative losses, and nonlinear coupling between velocity, magnetic fields, and thermodynamic variables. These nonlinear, scale-coupled interactions make it extremely difficult to isolate causal drivers analytically or via direct term-by-term decomposition. In contrast, IT tools, like NMI, allow empirical quantification of statistical dependencies, without assuming specific forms of the governing equations. As high-resolution chromospheric observations continue to improve, IT diagnostics offer a practical and scalable tool set for interpreting complex magnetic structures such as vortices, swirls, or flux ropes. Ultimately, these methods enable new ways to disentangle the drivers of plasma heating and energy transport in the solar atmosphere, and extend naturally to other astrophysical systems governed by turbulent, nonlinear dynamics.
@@ Acknowledgments
S.S.A.S. and V.F. are grateful to the Science and Technology Facilities Council (STFC) grants ST/V000977/1, ST/Y001532/1. S.S.A.S., I.B., V.F. are grateful to The Royal Society, International Exchanges Scheme, collaboration with Greece (IES/R1/221095). E.L.R. acknowledges the Brazilian agency CNPq under grant 306920/2020-4. S.S.A.S. and V.F. would like to thank the International Space Science Institute (ISSI) in Bern, Switzerland, for the hospitality provided to the members of the team on ‘Opening new avenues in identifying coherent structures and transport barriers in the magnetised solar plasma’. S.S.A.S. is grateful to Iñaki Esnaola for insightful discussions that contributed to this work.

### v2

@@ 2.5 Plasma heating
To identify the key drivers of temperature evolution within the vortex, we computed the NMI between plasma temperature and a selection of physical variables linked to heating processes or thermodynamic coupling. Figure 10 presents the temporal and vertical variation of these dependencies within the vortex core. The initial enhanced coupling between , and temperature is most pronounced during the vortex development and in the lower atmosphere, suggesting that the initial plasma cooling is driven by the jet transporting the vortex upwards and the initial flux tube expansion. As expected, plasma temperature initially shows strong statistical coupling with both density and pressure, consistent with the ideal gas law under quasi-static conditions. However, during the vortex phase, the coupling with plasma density significantly weakens above 1 Mm, indicating that the temperature evolution is no longer primarily governed by adiabatic compression or expansion. Instead, this indicates a breakdown of ideal thermodynamic behaviour due to non-adiabatic processes, such as current dissipation, wave heating, and magnetically induced flows. The reduced NMI between temperature and density/pressure reflects the growing influence of magnetic forces and energy injection mechanisms in the upper atmosphere.
@@
Although vortex-driven expansion initially leads to adiabatic cooling, this effect is short-lived. Within 10 time steps, the plasma temperature begins to rise, indicating that additional heating mechanisms rapidly become dominant. The sustained NMI between temperature and magnetic shear throughout the vortex lifetime suggests that current structures play a consistent role in modulating plasma heating. This coupling is particularly notable given that the magnetic shear-temperature association emerges primarily during the vortex phase, vanishing before and after. This timing supports the interpretation that currents, associated with twisted and evolving magnetic fields, contribute to localized heating within the vortex core. The temporal alignment between this temperature rise and elevated magnetic shear points toward current dissipation as a likely source of energy input. Although this dissipation is numerical in origin, because of the lack of explicit resistivity in the simulation, it still reflects the physical tendency for currents to form and dissipate in regions of strong shear. These results indicate that vortex-driven magnetic structuring channels energy into heat, particularly where plasma- is low (upper atmosphere)
3 Discussion
@@
To understand the vortex-driven dynamics in a simulated solar atmospheric plasma, we examined two large magnetic flux tubes: one that hosts a strong solar tornado with a coherent vortex flow and another characterized only by shear flows and transient swirls. The vortex flow first emerges in the upper photosphere and lower chromosphere. From there, it expands both upward and downward, forming a long-lived column of rotating plasma. It persists for approximately 5-10 minutes, with a horizontal radius of over 1 Mm in the upper atmosphere. Our analysis focuses on its evolution from the photosphere through the upper atmosphere, within the height range of 0-4 Mm. Throughout the full height of the vortex, the plasma- steadily decreases as the structure develops, reflecting a growing dominance of magnetic forces over gas pressure. Interestingly, the magnetic field intensifies at the vortex onset and remains strong even after the large-scale structure dissolves, pointing to lasting magnetic reconfiguration as reported by Diaz_2024. Aside from a brief initial cooling, the vortex phase is marked by a persistent temperature increase, highlighting the efficiency of heating mechanisms operating throughout its lifetime.
@@
Using information-theoretic diagnostics, we quantified how vortex evolution alters the behaviour of key MHD variables. SE captures changes in the spatial complexity of individual fields, whereas NMI measures how relationships between different variables strengthen or weaken over time and height. This difference in what each metric probes is the sense in which they are complementary: SE describes the intrinsic structural evolution of a single variable, while NMI characterizes how that variable co-varies with others. The magnetic-field variables illustrate this complementarity clearly. The evolution of SE shows that the entropy of the vertical magnetic field increases during vortex onset in the upper atmosphere, indicating the development of stronger spatial gradients and more irregular distributions. As the vortex matures, entropy decreases, reflecting the emergence of a more ordered, collimated structure. In contrast, the magnetic-shear entropy starts low, rises sharply as currents intensify during vortex development, and falls again during the decay phase, consistent with current dissipation. Cross-comparison with NMI maps reveals that when magnetic-shear entropy peaks, temperature becomes strongly coupled to shear, indicating that thermal evolution is linked to the buildup and subsequent dissipation of current-related complexity.
@@
The thermodynamic variables also show clear stratification. In the lower atmosphere, SE indicates low entropy in density but high entropy in pressure, reflecting relatively uniform mass loading alongside strong compressive variability. NMI confirms that temperature remains persistently coupled to horizontal compression () throughout the vortex lifetime. At the same time, temperature also shows strong coupling with vorticity, which itself contributes to the velocity gradient. This dual dependence suggests that both compressive processes and viscous dissipation contribute to heating in the lower layers. In the upper atmosphere, the entropy patterns reverse: density entropy rises (stronger gradients and irregular distributions) while pressure entropy decreases (more ordered stratification). The coupling of temperature and density diminishes considerably in the upper part of the domain, indicating a departure from the ideal gas law. NMI reveals that temperature coupling with vorticity persists here, implying a continuing role for viscous heating, but now temperature also couples strongly to magnetic shear. This indicates that, higher up, magnetic structuring and current dissipation become additional drivers of thermal evolution alongside viscous effects. These findings confirm that vortices intensify velocity and magnetic field gradients, establishing conditions conducive to energy dissipation, consistent with previous studies (Yadav_2021; Silva_2024b; Silva_2024a). This behaviour is consistent with previous studies (Yadav_2021; Silva_2024b; Silva_2024a). Our analysis extends this picture by showing that plasma temperature is statistically coupled to these gradients, demonstrating that the dissipated energy is effectively converted into heating within the vortex. This means that regions where the vortex generates strong shear or currents correspond to sites of effective energy release.
@@
The coupling is not one-sided: plasma variables also influence vorticity. Elevated temperature can enhance pressure gradients, which in turn drive flows that sustain or modify vortical motions. In the lower atmosphere, where pressure entropy is high, pressure gradients are likely to reinforce rotational shear and help maintain vortex coherence. In the upper atmosphere, strong magnetic shear can exert Lorentz forces that realign flows, thereby shaping or even destabilizing the vortex. The observed NMI links, therefore, suggest a bidirectional relationship: vorticity drives gradients that heat the plasma, while thermal and magnetic structuring provide feedback that regulates the strength, persistence, and eventual decay of the vortex as reported by Shelyag_2012; Silva_2024b.
4 Conclusions
We applied SE and NMI to a long-lived vortex in a Bifrost simulation to quantify how it reshapes its local environment. Our analysis leads to the following key conclusions:
The vortex transforms the local dynamical regime. It suppresses the p-mode-like, pressure-dominated behaviour present in a neighbouring flux tube and replaces it with a magnetically controlled structure.
Magnetic complexity rises and falls with the vortex life cycle. SE peaks during vortex growth, consistent with current formation and energy buildup, and decreases as the field becomes more ordered.
Heating pathways change with height. Temperature responds mainly to compressional and expansion motions in the photosphere and lower chromosphere, while viscous and current-driven processes dominate in the upper atmosphere.
The plasma departs from adiabatic behaviour. Temperature-density coupling weakens during the vortex phase, indicating that dissipative processes increasingly govern the thermal evolution.
The strong couplings are generated by the vortex itself. A nearby non-vortical flux tube shows only weak and uniform NMI patterns, confirming that the enhanced links in the vortex region arise dynamically rather than from static co-location.
Information theory reveals interactions that traditional diagnostics miss. By distinguishing genuine dynamical coupling from mere spatial coincidence, NMI provides a quantitative view of how flows, fields, and thermodynamic variables interact during the vortex life cycle.
These results show that solar vortices are not passive features of turbulent flows. They actively reorganize the plasma, reshape magnetic fields, and drive localized, height-dependent heating. Information-theoretic tools also offer practical advantages for studying solar plasmas. NMI and SE provide a robust way to probe nonlinear interactions without relying on restrictive assumptions about the MHD equations. In the chromosphere, effects such as partial ionization, radiative losses, anisotropic conduction, and strong magnetic structuring limit the validity of the usual single-fluid, ideal-gas approach. This makes analytical separation of heating terms very difficult. The governing equations are highly nonlinear, coupled, and spread across multiple scales, which complicates direct term-by-term decomposition. In contrast, information-theoretic diagnostics quantify statistical dependencies directly from the data, allowing us to assess dynamical coupling and potential causal pathways even when the governing equations cannot be cleanly separated. As high-resolution chromospheric observations continue to improve, these methods offer a scalable, model-agnostic framework for interpreting complex magnetic structures such as vortices, swirls, and flux ropes, and they extend naturally to other turbulent and nonlinear astrophysical systems.
@@ Acknowledgments
S.S.A.S. and V.F. are grateful to the Science and Technology Facilities Council (STFC) grants ST/V000977/1, ST/Y001532/1. V.F., S.S.A.S., I.B. and G.V. are thankful to The Royal Society, International Exchanges Scheme, collaboration with Greece (IES/R1/221095) and The Royal Society, IEC/R3/233017 - International Exchanges 2023 Cost Share (NSTC), collaboration with Taiwan. E.L.R. acknowledges the Brazilian agency CNPq under grant 306920/2020-4. S.S.A.S. and V.F. would like to thank the International Space Science Institute (ISSI) in Bern, Switzerland, for the hospitality provided to the members of the team on ‘Opening new avenues in identifying coherent structures and transport barriers in the magnetised solar plasma’. S.S.A.S. is grateful to Iñaki Esnaola for insightful discussions that contributed to this work.